rvc (rvc):

please help me Using Lagrange's method,find minimum and maximum distance from point (1,2,2) to the sphere x^2+y^2+z^2=36.

OpenStudy (baru):

which part you having a problem with? or having trouble getting started?

OpenStudy (baru):

the part about the sphere represents the "constraint". You have to derive the equation that need to be min/maximised let A (x,y.z) be an arbitrary point in space, then distance between P and X is given by \[\sqrt{(x-1)^2+(y-2)^2+(z-2)^2}\]

OpenStudy (baru):

another helpful piece of info is that you can square the function for distance and then find max/min

OpenStudy (baru):

*distance between P and A

rvc (rvc):

hmm im bad at partial i require a lot of help

rvc (rvc):

@SolomonZelman @aryandecoolest

rvc (rvc):

i need to learn it pls show me the steps

rvc (rvc):

@sweetburger @Frostbite @shamil98 @ganeshie8 @AravindG @Nishant_Garg

rvc (rvc):

@zepdrix @Loser66 @iambatman pls help :'(

OpenStudy (baru):

how far have you understood lagranges method?

rvc (rvc):

hmm we have two functions F(x,y)=f(x,y)+g(x,y) sorry for late reply my laptop got hanged

OpenStudy (baru):

how did you get that?

OpenStudy (baru):

to apply lagranges method you need two unconnected functions w=f(x,y,z) and g(x,y,z)=c (constraint)

rvc (rvc):

okay

OpenStudy (baru):

the function of the sphere represents g(x,y,z) = c (c is 36)

rvc (rvc):

so in this case our sphere will be a constriant

OpenStudy (baru):

yep

rvc (rvc):

yess

OpenStudy (baru):

lagranges method states: when you have a function f and a constraint 'g' then, at max/min points \[\nabla f = \lambda \nabla g\] where 'lambda' is an arbitrary scalar quantity

rvc (rvc):

oh i dont know the meaning of that symbol

rvc (rvc):

sorry i m really bad at partial

OpenStudy (baru):

\[\nabla f =< \partial f/\partial x , \partial f/\partial y, \partial f/\partial z>\]

rvc (rvc):

oh okay

OpenStudy (baru):

you know how to do partial differentiation? you just have to treat all the other variables as constants

rvc (rvc):

yes i know that

OpenStudy (baru):

try finding \[\nabla g\]

rvc (rvc):

\[\tt g_x =2x\\ g_y=2y\\ g_z=2z\]

OpenStudy (baru):

yep :)

rvc (rvc):

yay

OpenStudy (baru):

you understood the equation for 'f' ?

OpenStudy (baru):

the one at the top of this thread, under the square root sign.

rvc (rvc):

yea it is the distance equation

rvc (rvc):

what is P there? centre of sphere?

OpenStudy (baru):

no, P is a random point somewhere inside the sphere, you need to find which point on the sphere is closest and farthest from P

OpenStudy (baru):

you understood that part where we said we can square the distance function?

rvc (rvc):

yep d^2=-----

rvc (rvc):

so from point (1,2,2) we are finding a point which is closest to it and far from it,correct?

OpenStudy (baru):

yep

rvc (rvc):

yes...

OpenStudy (baru):

so, you have understood that the distance function is the "f(x,y,z)" ?

rvc (rvc):

hmmm.. tbh no :(

rvc (rvc):

well for sure we have a function with constant so the other function has to be the f(x,y,z) function

OpenStudy (baru):

f(x,y,z) represents the function that needs to minimized/maximized. the question is asking you to minimize or maximize the distance between P and a point on the circle right.

rvc (rvc):

yes @baru

OpenStudy (anonymous):

Using Lagrange's method,find minimum and maximum distance from point (1,2,2) to the sphere x^2+y^2+z^2=36. so lets write our objective as the squared distance from \((1,2,2)\): $$r=(x-1)^2+(y-2)^2+(z-2)^2$$ now we want to extremize \(r\) with respect to the constraint \(x^2+y^2+z^2=36\); we can add a penalty for violating this constraint using a Lagrange multiplier \(\lambda\): $$r=(x-1)^2+(y-2)^2+(z-2)^2-\lambda(36-x^2+y^2+z^2)$$

OpenStudy (anonymous):

now extremize that function: $$\frac{\partial r}{\partial x}=2(x-1)+2\lambda x\\\frac{\partial r}{\partial y}=2(y-2)-2\lambda y\\\frac{\partial r}{\partial z}=2(z-2)-2\lambda z$$

OpenStudy (anonymous):

oops, it should read \(36-x^2-y^2-z^2\) so the sign of the \(\lambda\) terms in the partial derivatives should be positive

OpenStudy (anonymous):

so $$x=1-\lambda\\y=2-\lambda\\z=2-\lambda$$ when \(r_x=r_y=r_z=0\)

OpenStudy (anonymous):

and our constraint becomes $$(1-\lambda)^2+(2-\lambda)^2+(2-\lambda)^2=36\\9-10\lambda+3\lambda^2=36\\3\lambda^2-10\lambda-27=0$$

OpenStudy (anonymous):

so: $$\lambda=\frac{10\pm\sqrt{100+4(81)}}{6}=\frac{5\pm\sqrt{106}}3$$

OpenStudy (anonymous):

plug this in for \(x,y,z\) for the points with minimum and maximum distance, respectively

OpenStudy (baru):

@rvc followed?

rvc (rvc):

i would love to hear your explainah]tion too @baru

rvc (rvc):

explaination*

OpenStudy (baru):

are you clear on part about "constraint" and the function "f"?

rvc (rvc):

yepyep

OpenStudy (baru):

so rather than suffer with an equation under the square root, we convinently choose square of distance for "f". can you find \[\nabla f\]?

rvc (rvc):

\[\rm d^2=(x-1)^2+(y-2)^2+(z-2)^2\]

rvc (rvc):

please reply to my message

OpenStudy (baru):

now f = d^2 find\[\nabla f\]

rvc (rvc):

\[\rm f_x=2(x-1)\\ f_y=2(y-1)\\ f_z=2(z-1)\]

OpenStudy (baru):

yep

OpenStudy (baru):

now try \[\nabla f = \lambda \nabla g\]

OpenStudy (baru):

the first equation you get would be \[2(x-1)= \lambda 2x\]

OpenStudy (baru):

can you do the other two?

rvc (rvc):

\[\tt 2(y-1)=2y\\ 2(z-1)=2z\]

rvc (rvc):

oops i forgot the \(\lambda\)

OpenStudy (baru):

:) ya

rvc (rvc):

yup :)

rvc (rvc):

well sorry to message you. I thought you wouldn't mind me :) its fine thanks for all the help though

OpenStudy (baru):

message? you mean pm? i didnt get any

rvc (rvc):

i did message you 3-4 times

OpenStudy (baru):

oh..lol. didnt get any notifications

rvc (rvc):

now can u please relpyto my messages Sir

OpenStudy (baru):

yep :)

OpenStudy (baru):

@rvc are we done with this question? you know how to proceed?

rvc (rvc):

further i have to solve for lambda right

OpenStudy (baru):

thats upto you, in the end you just need to find values for x,y and z which satisfy those three equations and the constraint equation 'g=c'

rvc (rvc):

oh i need more help

OpenStudy (baru):

the best way to do this would be to get rid of lambda \[2(x-1)/2x = \lambda\] \[2(y-2)/2y= \lambda\] \[2(z-2)/2z=\lambda\]

OpenStudy (baru):

now see how all three equations are equal to lambda and thus equal to each other

rvc (rvc):

yesss